Simplification of Context-Free Grammars and Normal Forms

4. Transform the grammar with productions

S → baAB,

A → bAB|λ,

B → BAa |A| λ

into CNF and GNF


 5. Convert the grammar

S → AB|aB,

A → abb|λ,

B → bbA

into CNF and GNF.
Answer #1

Following is the procedure to transform:-

1. We will remove λ-productions: –

Removing A → λ: S → baAA|baA|ba, A → bAB|bB, B → BAa|A|Ba.

Removing B → λ: S → baAA|baA|baA|ba, A → bAB|bB|bA|b, B → BAa|A|Ba|Aa|a.

2. We will remove unit-production:-

B → A: S → baAA|baA|ba, A → bAB|bB|bA|b, B → BAa|bAB|bB|bA|b|Ba|Aa|a.

3. Now convert the grammar into Chomsky normal form: –

a) We will introduce new variables Sa for each a ∈ T and Sb for each b ∈ T:

S → SbSaAA|SbSaA|SbSa,

A → SbAB| SbB| SbA| Sb,

B → BASa|SbAB|SbB|SbA|Sb|B|ASa|Sa,

Sa → a,

Sb→ b.

b) Now we are introducing additional variables to get the first two productions into normal form and we get the final result

S → SbU|SbX |SbSa,

A → SbY |SbB|SbA|Sb,

B → BZ|SbY |SbB|SbA|Sb|BSa|ASa|Sa,

U → SaV,

V → AA,

X →Sa A,

Y → AB,

Z → ASa,

Sa → a,

Sb → b.
Answer #2

Production given are as follows :Converting the grammar into Chomsky normal form: \(-\) Introduce new variables \(\mathrm{S}_{\mathrm{a}}\) for each \(\mathrm{a} \in \mathrm{T}\) and \(\mathrm{S}_{\mathrm{b}}\) for each \(\mathrm{b} \in \mathrm{T}\) :

\(\mathrm{S} \rightarrow \mathrm{S}_{\mathrm{b}} \mathrm{S}_{\mathrm{a}} \mathrm{A} \mathrm{B}\left|\mathrm{S}_{\mathrm{b}} \mathrm{S}_{\mathrm{a}} \mathrm{B}\right| \mathrm{S}_{\mathrm{b}} \mathrm{S}_{\mathrm{a}} \mathrm{A} \mid \mathrm{S}_{\mathrm{b}} \mathrm{S}_{\mathrm{a}}\)

\(\mathrm{A} \rightarrow \mathrm{S}_{\mathrm{b}} \mathrm{AB}\left|\mathrm{S}_{\mathrm{b}} \mathrm{B}\right| \mathrm{S}_{\mathrm{b}} \mathrm{A} \mid \mathrm{S}_{\mathrm{b}}\)

\(\mathrm{B} \rightarrow \mathrm{BAS}_{\mathrm{a}}\left|\mathrm{S}_{\mathrm{b}} \mathrm{AB}\right| \mathrm{S}_{\mathrm{b}} \mathrm{B}\left|\mathrm{S}_{\mathrm{b}} \mathrm{A}\right| \mathrm{S}_{\mathrm{b}}\left|\mathrm{BS}_{\mathrm{a}}\right| \mathrm{AS}_{\mathrm{a}} \mid \mathrm{S}_{\mathrm{a}}\)

\(\mathrm{S}_{\mathrm{a}} \rightarrow \mathrm{a}\)

\(\mathrm{S}_{\mathrm{b}} \rightarrow \mathrm{b}\)

Introducing more variables to get the productions into normal form and the final result are as follows:

\(\mathrm{S} \rightarrow \mathrm{S}_{\mathrm{b}} \mathrm{U}\left|\mathrm{S}_{\mathrm{b}} \mathrm{X}\right| \mathrm{S}_{\mathrm{b}} \mathrm{Y} \mid \mathrm{S}_{\mathrm{b}} \mathrm{S}_{\mathrm{a}}\)

\(\mathrm{A} \rightarrow \mathrm{S}_{\mathrm{b}} \mathrm{V}\left|\mathrm{S}_{\mathrm{b}} \mathrm{B}\right| \mathrm{S}_{\mathrm{b}} \mathrm{A} \mid \mathrm{S}_{\mathrm{b}}\)

\(\mathrm{B} \rightarrow \mathrm{BZ}\left|\mathrm{S}_{\mathrm{b}} \mathrm{V}\right| \mathrm{S}_{\mathrm{b}} \mathrm{B}\left|\mathrm{S}_{\mathrm{b}} \mathrm{A}\right| \mathrm{S}_{\mathrm{b}}\left|\mathrm{BS}_{\mathrm{a}}\right| \mathrm{AS}_{\mathrm{a}} \mid \mathrm{S}_{\mathrm{a}}\)

\(\mathrm{U} \rightarrow \mathrm{S}_{\mathrm{a}} \mathrm{V}\)

\(\mathrm{V} \rightarrow \mathrm{AB}\)

\(\mathrm{X} \rightarrow \mathrm{S}_{\mathrm{a}} \mathrm{B}\)

\(\mathrm{Y} \rightarrow \mathrm{S}_{\mathrm{a}} \mathrm{A}\)

\(\mathrm{Z} \rightarrow \mathrm{AS}_{\mathrm{a}}\)

\(\mathrm{S}_{\mathrm{a}} \rightarrow \mathrm{a}\)

\(\mathrm{S}_{\mathrm{b}} \rightarrow \mathrm{b}\)


\(\mathrm{S} \rightarrow \mathrm{baAB}\)

\(\mathrm{A} \rightarrow \mathrm{bAB} \mid \lambda\)

\(\mathrm{B} \rightarrow \mathrm{BAa}|\mathrm{A}| \lambda\)

Removing the \(\lambda\) -productions:

First removing \(A \rightarrow \lambda:\)

\(\mathrm{S} \rightarrow \mathrm{baAB} \mid \mathrm{baB}\)

\(\mathrm{A} \rightarrow \mathrm{AB} \mid \mathrm{bB}\)

\(\mathrm{B} \rightarrow \mathrm{BAa}|\mathrm{A}| \lambda \mid \mathrm{Ba}\)

Then removing \(\mathrm{B} \rightarrow \lambda\)

\(\mathrm{S} \rightarrow \mathrm{baAB}|\mathrm{baB}| \mathrm{baA} \mid \mathrm{ba}\)

\(\mathrm{A} \rightarrow \mathrm{bAB}|\mathrm{bB}| \mathrm{bA} \mid \mathrm{b}\)

\(\mathrm{B} \rightarrow \mathrm{BAa}|\mathrm{A}| \mathrm{Ba}|\mathrm{Aa}| \mathrm{a}\)

Removing unit-production \(\mathrm{B} \rightarrow \mathrm{A}\) :

\(\mathrm{S} \rightarrow \mathrm{baAB}|\mathrm{baB}| \mathrm{baA} \mid \mathrm{ba}\)

\(\mathrm{A} \rightarrow \mathrm{bAB} / \mathrm{bB}|\mathrm{bA}| \mathrm{b}\)

\(\mathrm{B} \rightarrow \mathrm{BAa}|\mathrm{bAB}| \mathrm{bB}|\mathrm{bA}| \mathrm{b}|\mathrm{Ba}| \mathrm{Aa} \mid \mathrm{a}\)


answered by: Aiden